4. A model for a company's revenue from selling a software package is R = -2.5p2 + 500p, where p is the price in dollars of the software.
a. What price will maximize the revenue? Give a reason for your answer.
b. What is the maximum revenue? Give a reason for your answer.

Found 2 solutions by TimothyLamb, ewatrrr:
Answer by TimothyLamb(4379)

R = -2.5p2 + 500p
---
substitute p with x
set R = 0
-2.5xx + 500x = 0
---
the above quadratic equation is in standard form, with a=-2.5, b=500 and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-2.5 500 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has a maximum at = ( 100, 25000 )
---
answer:
price for maximum revenue = $100
this is the p (or x) coordinate of the equation's maximum
---
maximum revenue = $25,000
this is the R (or y) coordinate of the equation's maximum
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by ewatrrr(24785)

R = -2.5p2 + 500p
Note: 500/-(-2*2.5) = 100
R = -2.5(p -100)^2 + 25, 000 Parabola opening downward V(100,25,000)
.........
a) p = $100 will maximize the revenue. V(100,25000)max
b) max R = $25,000 . V(100,25000)max